A hardware manufacturer produces bolts used to assemble various machines. Assume that the diameter of bolts produced by this manufacturer has an unknown population mean 𝝁 and the standard deviation is 0.1 mm. suppose the average diameter of a simple random sample of 50 bolts is 5.11 mm. (a) Calculate the margin of error of a 95% confidence interval for 𝝁.


1
Expert's answer
2022-05-24T10:11:10-0400

The critical value for "\alpha = 0.05" is "z_c = z_{1-\alpha/2} = 1.96."



"ME=z_c\times\dfrac{\sigma}{\sqrt{n}}=1.96\times\dfrac{0.1}{\sqrt{50}}""=0.0277"